### v1

# LoKI: Low-damage Knowledge Implanting of Large Language Models

Abstract
Fine-tuning adapts pretrained models for specific tasks but poses the risk of catastrophic forgetting (CF), where critical knowledge from pre-training is overwritten. Current Parameter-Efficient Fine-Tuning (PEFT) methods for Large Language Models (LLMs), while efficient, often sacrifice general capabilities. To address the issue of CF in a general-purpose PEFT framework, we propose Low-damage Knowledge Implanting (LoKI), a PEFT technique that is based on a mechanistic understanding of how knowledge is stored in transformer architectures. In two real-world scenarios, LoKI demonstrates task-specific performance that is comparable to or even surpasses that of full fine-tuning and LoRA-based methods across various model types, while significantly better preserving general capabilities. Our work connects mechanistic insights into LLM knowledge storage with practical fine-tuning objectives, achieving state-of-the-art trade-offs between task specialization and the preservation of general capabilities. Our implementation is publicly available as ready-to-use code [^1] https://github.com/Nexround/LoKI.
1 Introduction
Can we develop a Parameter-Efficient Fine-Tuning (PEFT) [prottasha_peft_2025, xin_parameter-efficient_2025, han_parameter-efficient_2024] method that can make the Large Language Models (LLMs) [radford2018improving, brown_language_2020] work smoothly on the target task after fine-tuning, while preventing catastrophic forgetting (CF) [DBLP:conf/iclr/KothaSR24, luo_empirical_2025]?
While PEFT techniques such as LoRA [hu2022lora] have significantly reduced the cost of adapting LLMs to downstream tasks, they often fail to preserve the model’s general capabilities. A key challenge is CF—the tendency of the model to lose general knowledge acquired during pretraining [jiang_how_2020, petroni_language_2019] when fine-tuned on specific tasks.
Recent mechanistic studies reveal that factual knowledge in LLMs associations Feed-Forward Network (FFN) layers in the transformer blocks [vaswani_attention_2017, geva_transformer_2021-1, DBLP:conf/iclr/MengSABB23]. Conventional fine-tuning approaches, however, apply indiscriminate updates across attention layers or FFN blocks, oblivious to these crucial knowledge-storing weights in general tasks. Such unguarded optimization may perturb crucial memory traces [cohen_crawling_2023, petroni_language_2019], leading to irreversible knowledge loss and degraded generalization [li_revisiting_2024, DBLP:conf/acl/HuangCWYLSYS24] - the dilemma of modern LLM adaptation.
We propose Low-damage Knowledge Implantation (LoKI), a PEFT framework that achieves state-of-the-art performance in CF mitigation while maintaining competitive task adaptation. Our approach is grounded in two key insights:
FFN layers as primary knowledge hubs. Recent studies on knowledge localization and model editing have revealed that factual and semantic knowledge in transformer models is primarily stored within the FFN layers [DBLP:conf/aaai/Li0SYMY24, DBLP:conf/iclr/MengSABB23, DBLP:conf/iclr/TanZF24, DBLP:journals/corr/abs-2310-10322, DBLP:conf/iclr/FangJWMSW0C25]. Individual weight vectors and neurons are found to express distinct facts [DBLP:conf/acl/DaiDHSCW22, geva_transformer_2022-1]. Furthermore, it has been confirmed that specific expressions of knowledge in transformer-based language models can be altered by editing neuron activations or adjusting layer weights in FFNs [DBLP:conf/acl/DaiDHSCW22, meng_locating_2022, de_cao_editing_2021-1, DBLP:conf/iclr/MengSABB23].
Redundancy in pretrained LMs. Extensive work on pruning and sparsity in language models [DBLP:conf/iclr/HeZMBN22, DBLP:conf/iclr/FrankleC19, DBLP:conf/icml/FrantarA23, DBLP:conf/nips/MichelLN19, voita_analyzing_2019] observed that LLMs harbor a substantial number of "dormant" or low-impact weights, especially in the FFN layers [DBLP:conf/iclr/KobayashiKYI24, DBLP:conf/nips/Sanh0R20, DBLP:conf/acl/ZhangL00S022]. We proved that these low-impact weights in FFN layers, located in our experiments, create safe "implant slots" where new, task-specific knowledge can be implanted without a significant impact on the general performance that LMs obtained during pre-training.
By combining these two principles—allocating updates to carefully selected FFN weights—LoKI redefines the paradigm for sustainable LLM customization, ensuring models evolve without losing their essence. In summary, our main contributions are:
  • We introduce Low-damage Knowledge Implantation (LoKI), a PEFT method that simultaneously (a) adapts to downstream tasks and (b) mitigates CF in pre-trained LMs.
  • We propose Knowledge Vector Attribution (KVA), a technique to quantify the contribution of individual knowledge vectors (defined in Section 2) to the model’s stored representations. Through ablation studies, we demonstrate that KVA effectively identifies safe "implant slots" for task-specific updates.
  • Building on an understanding of the hierarchical organization of knowledge in transformers, we develop the Layer-Balanced Strategy for allocating trainable weights. Our experiments show that this strategy is critical to LoKI’s ability to preserve pre-trained capabilities while learning new tasks.
2 Knowledge in Transformer-Based Language Models
In this section, we will begin by briefly introducing the FFN layer in transformers [vaswani_attention_2017] and provide some definitions to support our discussion of the method. Following that, we will review relevant research on knowledge storage in transformer-based language models, which will establish a foundation for our proposed method.
Modern language models are built using a series of transformer layers. Each standard transformer layer contains two main components: (1) multi-head self-attention (MHSA) [vaswani_attention_2017] and (2) a position-wise feed-forward network. In this study, we focus on the FFN component, as recent research has indicated that FFNs serve as knowledge repositories within language models.
Feed-Forward Network. Consider a FFN without bias, given an input vector $x\in\mathbb{R}^{d_{\mathrm{model}}}$, the FFN applies two linear projections with a nonlinearity in between:
$\mathrm{FFN}(x)\;=W_{\mathrm{down}}\;\sigma\bigl{(}W_{\mathrm{up}}\,x\bigr{)}$ (1)
where $W_{\mathrm{up}}\in\mathbb{R}^{d_{\mathrm{ff}}\times d_{\mathrm{model}}}$ is the "up-projection" matrix, $W_{\mathrm{down}}\in\mathbb{R}^{d_{\mathrm{model}}\times d_{\mathrm{ff}}}$ is the "down-projection" matrix, $\sigma(\cdot)$ is an element-wise nonlinearity.
Definitions. To facilitate discussion, define the intermediate activation and the final output node:
$a=\sigma(W_{\mathrm{up}}\,x)\in\mathbb{R}^{d_{\mathrm{ff}}},\quad y=W_{\mathrm{down}}\,a\in\mathbb{R}^{d_{\mathrm{model}}}.$ (2)
Denote the $j$th output coordinate by $y_{j}=v_{j}^{\top}a,\quad v_{j}=W_{\mathrm{down},j\cdot}\in\mathbb{R}^{d_{\mathrm{ff}}}.$ We call $y_{j}$ a Knowledge Output Node and its vector $v_{j}$ a Knowledge Vector. Given the one-to-one correspondence between these two entities, we will use these terms interchangeably throughout the following sections.
FFNs as key-value memories.
FFN naturally implements a key-value memory [geva_transformer_2021-1]: each column of $W_{\mathrm{up}}$ serves as a key that detects specific input patterns, and each row of $W_{\mathrm{down}}$ acts as a value corresponding to a distinct concept in the lexical space [geva_transformer_2021-1, geva_transformer_2022-1]. This mechanism has also been validated from the perspective of gradient analysis [katz_backward_2024]. Furthermore, Geva et al. [geva_transformer_2022-1] have revealed that FFN output updates can be decomposed into multiple sub-updates induced by value vectors (knowledge vectors) in $W_{\mathrm{down}}$. Based on this insight, our focus will be exclusively on updating the parameters within $W_{\mathrm{down}}$.
Hierarchical structure of layers.
While single-layer FFNs reveal this memory behavior, deeper inspection shows a hierarchy: lower FFN layers capture surface-level patterns, whereas upper layers encode higher-level semantics [geva_dissecting_2023, DBLP:conf/acl/DaiDHSCW22, DBLP:conf/iclr/TanZF24, katz_backward_2024]. The final output distribution of transformer-based language models is gradually constructed in a bottom-up fashion [DBLP:conf/aaai/SunPNJ25, tenney_bert_2019, wallat_bertnesia_2020].
However, prior studies have demonstrated that LLMs contain a substantial number of parameters whose removal has minimal impact on the model’s performance on its primary tasks [DBLP:conf/iclr/FrankleC19, DBLP:conf/iclr/LasbyGENI24, DBLP:conf/nips/MichelLN19], particularly in FFN layers [DBLP:conf/acl/ZhangL00S022, sanh_movement_2020, DBLP:conf/icml/FrantarA23]. Based on the previous research results mentioned above, we hypothesize that these low-impact parameters in memory components can serve as "safe zones" that can be repurposed to encode new task-specific knowledge without excessively degrading the model’s original capabilities. Encouragingly, our experimental results support this hypothesis.
3 Low-damage Knowledge Implanting
Figure 1: Schematic illustration of LoKI.
We introduce Low-damage Knowledge Implanting (LoKI), a three-stage framework—analyzing, selecting, and implanting—designed to edit transformer models with minimal disruption, thereby mitigating catastrophic forgetting (see Figure 1). Below, we briefly outline each stage before providing a detailed explanation:
  1. 1.
    Analyzing: Evaluate the contribution of each knowledge vector to general tasks.
  2. 2.
    Selecting: Choose trainable knowledge vectors within each FFN based on the analysis results.
  3. 3.
    Implanting: Train the chosen vectors to incorporate task-specific knowledge.
3.1 Analysing
We start this subsection by introducing Knowledge Vector Attribution (KVA), an approach inspired by [DBLP:conf/aaai/Hao0W021, DBLP:conf/acl/DaiDHSCW22]. Next, we will illustrate how KVA is applied in LoKI.
3.1.1 Knowledge Vector Attribution
KVA is a computational approach based on Integrated Gradients (IG) [DBLP:conf/icml/SundararajanTY17] that measures the contribution of knowledge vectors to specific output logits. To start with, we present the general formulation of IG:
[…]
where $F$ is the network function, $x^{\prime}$ is a baseline (we use $x^{\prime}=\mathbf{0}$), and $\alpha$ parameterizes the interpolation path. IG attributes predictions to input features by integrating gradients along the straight-line path from baseline $x^{\prime}$ to input $x$. To trace knowledge flow through all layers of a transformer on input sequence $\mathbf{x}$, let $\mathcal{L}$ denote the target logit, $h_{l-1}$ represent the hidden state from the previous layer. At layer $l$, the FFN first computes pre-activations
$z_{l}=W_{\mathrm{up}}^{(l)}\,h_{l-1},\quad\mathbf{u}_{l}=\sigma(z_{l}),\quad\mathbf{z}_{l,j}=\mathbf{u}_{l,j}\,W_{\mathrm{down},j\cdot}^{(l)}.$ (4)
We define the layer-wise, path-integrated attribution of node $j$ by
[…]
which aggregates the total gradient flowing through knowledge vector $j$ as its contribution is scaled from zero up to its actual activation. We approximate the above integral via Riemann approximation with $m$ equally-spaced steps (we set the step $m=7$ for any model):
$A_{l,j}(\mathbf{x})\approx\frac{1}{m}\sum_{k=1}^{m}\frac{\partial\mathcal{L}\bigl{(}\tfrac{k}{m}\,\mathbf{z}_{l,j}\bigr{)}}{\partial\mathbf{z}_{l,j}}.$ (6)
During each sample, KVA computes and stores $A_{l,j}(\mathbf{x})$ for every layer $l$ and every knowledge output node $j$. This yields a complete attribution log that can be analyzed post hoc to assess the contribution of each knowledge vector during inference.
3.1.2 Application to MMLU
We apply KVA on the Massive Multitask Language Understanding (MMLU) benchmark [DBLP:conf/aaai/KemkerMAHK18] to measure individual knowledge vector contributions to the pretrained model’s performance on general tasks. MMLU is ideal because:
  • it has 57 diverse subjects (STEM, humanities, professional, etc.), which covers a broad spectrum of general-knowledge tasks, ensuring extracted vectors are not domain-specialized;
  • its standardized prompts and evaluation make attributions comparable across tasks;
  • high performance on MMLU correlates with real-world language understanding, so protecting original performance is measurable.
We calculate IG scores of each FFN’s knowledge output nodes using the top 50 examples from each subject in the MMLU benchmark. We denote the recorded IG scores as $\mathbf{IG}\in\mathbb{R}^{N\times L\times D}$, where $N$ represents the number of examples, $L$ denotes the number of layers in the model, and $D$ indicates the dimensionality of the knowledge output nodes. We conducted this process using an A100 GPU for Llama3.1-8B-Instruct [DBLP:journals/corr/abs-2407-21783]. The average time required for each sample is 16.14 seconds. It’s important to note that the KVA needs to be executed only once for each model, regardless of the downstream tasks. In other words, this computational overhead occurs just once per model. Implementation details and further discussions can be found in Appendix A.
To determine if the attribution scores accurately represent the importance of individual knowledge vectors to overall task performance, we performed ablation experiments and presented the results in Section 5.1.
3.2 Selecting
As shown in Figure 2, the evaluation results of KVA reveal significant heterogeneity in the distribution of knowledge output nodes across different layers, regardless of their contribution level. Notably, both high-contribution and low-contribution nodes tend to be densely concentrated in a few middle layers and the top layers. This uneven distribution phenomenon among nodes with high contributions aligns with the findings of Meng et al. [meng_locating_2022, DBLP:conf/iclr/MengSABB23] and Dai et al. [dai_knowledge_2022-1]. Our work further observed the phenomenon of uneven distribution of low contribution nodes.
Given the established relationships between the layers within transformers [sun2025transformer, geva_transformer_2022-1], focusing weight updates on a few layers for downstream task adaptation may disrupt this hierarchical relationship. Our experiments revealed that this uneven parameter allocation will lead to a more pronounced phenomenon of catastrophic forgetting (see Section 5.2 for details). Additionally, Hase et al. [DBLP:conf/nips/HaseBKG23] have found a significant challenge: even when factual knowledge resides in specific layers, such as the mid-layer FFNs, modifying weights in distant layers, especially the earlier ones, can effectively "override" the information flow downstream.
Figure 2: Heatmaps of top $5\%$ KVA results across all 32 FFN layers of Llama3.1-8B-Instruct. The upper (red-tinted) map highlights the distribution of high-contribution node positions, while the lower (blue-tinted) map highlights the distribution of low-contribution node positions. Color intensity (log-scale) indicates the density of nodes in each category, with darker colors representing higher density. Heatmaps of other models can be found in Appendix B.
3.2.1 Layer-Balanced Strategy
Based on the earlier considerations, we propose the Layer-Balanced Strategy, which is a strategy that determines the trainable parameter positions under the constraint of allocating the same number of parameters to each layer. This strategy aims to (a) ensure that newly implanted knowledge conforms to the hierarchical relationship of the transformer model’s knowledge structure, and (b) avoid disproportionately disturbing its inherent knowledge hierarchy.
Let the model have $L$ layers, where only parameters within each layer’s $W_{\mathrm{down}}$ matrix are partially trainable, while all other parameters across all layers remain frozen. Each $W_{\mathrm{down}}$ in layer $l$ contains $D_{l}$ knowledge output nodes (with $D_{l}=D$ assumed identical across layers for simplicity). The hyperparameter $q\in(0,100)$ governs the percentage of trainable nodes selected from all $W_{\mathrm{down}}$ matrices. Given $N$ inference samples, we denote the integrated gradients for node $i$ in layer $l$ on sample $t$ as:
$\mathbf{IG}_{l,i}^{(t)},\quad i=1,\dots,D,\;t=1,\dots,N$
The implementation proceeds as follows:
  1. 1.
    Quota Allocation: Calculate the total trainable slots:
    $T=\frac{q}{100}\cdot\sum_{l=1}^{L}D_{l}=\frac{q}{100}\cdot LD$ (7)
    Allocate equal quotas per layer: $k_{l}=\left\lfloor\frac{T}{L}\right\rfloor,\quad l=1,\dots,L.$
  2. 2.
    Per-Sample Normalization & Local Selection: For each sample $t$ and layer $l$, normalize gradients:
    $\hat{\mathbf{IG}}_{l,i}^{(t)}=\frac{\mathbf{IG}_{l,i}^{(t)}-\min_{j}\mathbf{IG}_{l,j}^{(t)}}{\max_{j}\mathbf{IG}_{l,j}^{(t)}-\min_{j}\mathbf{IG}_{l,j}^{(t)}}$ (8)
    Select $k_{l}$ nodes with the smallest normalized values:
    $S_{l}^{(t)}=\underset{i\in\{1,\dots,D\}}{\mathrm{argsort}\downarrow}\left(\hat{\mathbf{IG}}_{l,i}^{(t)}\right)[1:k_{l}]$ (9)
  3. 3.
    Frequency Aggregation: Tally selection frequencies across samples:
    $c_{l,i}=\sum_{t=1}^{N}\mathbb{I}\left(i\in S_{l}^{(t)}\right)$ (10)
  4. 4.
    Final Selection: For each layer $l$, select nodes with the highest frequencies:
    $\mathcal{S}_{l}=\underset{i\in\{1,\dots,D\}}{\mathrm{argsort}\uparrow}(c_{l,i})[1:k_{l}],\quad\mathcal{S}=\bigcup_{l=1}^{L}\mathcal{S}_{l}$ (11)
    Resulting in a balanced set $\mathcal{S}$ with $\sum_{l=1}^{L}|\mathcal{S}_{l}|\leq T$.
3.3 Implanting
Building upon the selected knowledge output nodes $\mathcal{S}=\bigcup_{l=1}^{L}\mathcal{S}_{l}$ from Section 3.2, we implement parameter-efficient fine-tuning through strategic decomposition of FFN layers. For each layer $l$’s down-projection matrix $W_{\mathrm{down}}^{(l)}$, we partition the parameters into two complementary subspaces:
$W_{\mathrm{down}}^{(l)}=\begin{bmatrix}W_{\mathcal{S}_{l}}\\[3.0pt] W_{\setminus\mathcal{S}_{l}}\end{bmatrix},$ (12)
where $W_{\mathcal{S}_{l}}\in\mathbb{R}^{|\mathcal{S}_{l}|\times d_{\mathrm{ff}}}$ contains the weights corresponding to our selected low-contribution knowledge output nodes (from layer $l$’s quota $\mathcal{S}_{l}$), and $W_{\setminus\mathcal{S}_{l}}$ represents the remaining parameters. During training, we:
  • Keep $W_{\setminus\mathcal{S}_{l}}$ frozen to preserve existing knowledge representations
  • Update only $W_{\mathcal{S}_{l}}$ actively to implant new knowledge
This decomposition retains the mathematical formulation of the original layer while restricting parameter updates to the chosen subspaces. Additionally, it transforms LoKI training into a module-wise process, enabling easy integration with existing training pipelines. Notably, this implementation allows for the incorporation of low-rank decomposition techniques [hu2022lora, DBLP:conf/icml/LiuWY0WCC24, DBLP:conf/iclr/ZhangCBH0CZ23] into LoKI. To verify the feasibility of this combination, we conducted an experiment in Section 4.2. For clarity, we refer to this implementation as LoKI Linear, and readers can find its PyTorch code in Appendix C.
4 Experiments
In this section, we detail the experiments conducted to evaluate the effectiveness of the LoKI. We focus on two tasks: adapting models on the LB Reranker Dataset and the ToolACE Function-Calling Dataset. These publicly available datasets were specifically selected due to their demonstrated practical utility in real-world applications and their ability to simulate realistic PEFT demand scenarios. Each experiment assesses the capability of LoKI to mitigate catastrophic forgetting while acquiring task-specific performance. Notably, while our experiments involved only two model types, the proposed method is readily applicable to other model architectures.
  1. 1.
    LB Reranker Dataset [lb_reranker]: This multilingual dataset consists of 2.28 million query-text pairs annotated with fine-grained relevance scores on a 1-7 scale, designed for training NLP-based retrieval models. It has an official full parameter fine-tuning model based on Qwen2.5-0.5B-Instruct [qwen_qwen25_2025]; we utilized the same base model on this dataset. The experiment setup for this task can be found in Appendix D.
  2. 2.
    ToolACE Function-Calling Dataset [liu_toolace_2024]: This dataset contains 26,507 distinct APIs across various domains, designed to enhance the function-calling capabilities of LLMs. This richly diverse and rigorously validated dataset underpins state-of-the-art performance in function-calling benchmarks [berkeley-function-calling-leaderboard], even for 8B-parameter models. Consistent with the official model, we fine-tuned Llama3.1-8B-Instruct [grattafiori_llama_2024] on this dataset. The details of the experiment setup can be found in Appendix E.
For both tasks, to systematically evaluate CF during fine-tuning, we employ six benchmarks to assess model capabilities across diverse domains: TriviaQA [DBLP:conf/acl/JoshiCWZ17] (world knowledge), GSM8K [cobbe2021gsm8k] (mathematical reasoning), HellaSwag [DBLP:conf/acl/ZellersHBFC19] and WinoGrande [DBLP:conf/aaai/SakaguchiBBC20] (commonsense understanding), HumanEval [chen2021evaluating] (code generation), and IFEval [zhou_instruction-following_2023] (instruction following). The information regarding our evaluation of these benchmarks is in Appendix F.
4.1 Task 1: LB Reranker Dataset
Table 1: Comparison with the full parameter fine-tuning model on BEIR.
Model MAP@1 (%) MAP@10 (%) Recall@1 (%) Recall@10 (%) NDCG @1(%) NDCG @10(%) P@1 (%) P@10 (%)
LoKI($q{=}5$) -3.1 -5.9 -3.1 -3.5 -2.2 -4.0 -1.8 -3.6
LoKI($q{=}10$) -2.3 -3.8 -2.3 -2.1 -0.3 -2.2 -0.5 -2.0
LoKI($q{=}20$) 0.2 -2.5 0.2 -1.5 0.7 -1.4 0.8 -1.5
LoKI($q{=}30$) 1.0 -1.0 1.0 -0.5 2.1 -0.3 2.5 -0.5
Qwen2.5 -78.1 -77.7 -78.1 -68.3 -78.7 -71.8 -77.5 -65.4
Table shows the percentage difference in performance compared to the full parameter fine-tuning model across standard retrieval metrics. Bold represents the top performance score in each column; underline represents the runner-up. Qwen2.5 represents Qwen2.5-0.5B-Instruct
The BEIR benchmark [thakur_beir_2021] is a standardized evaluation framework for testing the performance of information retrieval models across a wide range of diverse datasets and tasks. Experimental results on the BEIR benchmark (Table 14) demonstrate that our LoKI method achieves competitive performance compared to full parameter fine-tuning while requiring substantially fewer trainable parameters. Specifically, LoKI with $q{=}30$ surpasses the full-parameter LB-Reranker baseline by an average of 0.54% across all evaluation metrics. Notably, it demonstrates absolute improvements in enhanced precision for the top-ranked retrieval results.
The progressive performance improvement with increasing trainable parameters (from $q{=}5$ to $q{=}30$) reveals a clear parameter-performance tradeoff. Even with only $q{=}20$, LoKI already achieves positive performance gains in four metrics.
Table 2: Benchmark scores of models fine-tuned on LB Reranker Dataset.
Model #Params TriviaQA GSM8K HellaSwag WinoGrande HumanEval IFEval Avg(%, ↓)
Qwen2.5 NaN 24.37 39.65 30.98 44.20 26.83 33.39 NaN
LoKI($q{=}5$) 5.1M 20.53 38.36 30.52 40.09 24.39 34.12 6.12
LoKI($q{=}10$) 10.4M 20.53 37.98 21.53 47.04 24.39 33.39 8.86
LoKI($q{=}20$) 20.9M 19.77 36.77 33.16 47.28 27.44 33.27 1.70
LoKI($q{=}30$) 31.3M 20.10 37.60 38.39 43.96 26.83 32.25 0.46
LB-Reranker 494M 4.65 2.65 0.00 0.00 0.00 23.20 84.13
LB-Reranker refers to the full parameter fine-tuning model released by the dataset provider. Avg denotes the average performance degradation percentage of each indicator compared to the original model indicator.
As shown in Table 2, LoKI models of all configurations exhibit significantly less performance degradation across all metrics compared to the full parameter fine-tuned model. Interestingly, when $q{=}30$, LoKI not only achieved the best averaged performance on the BEIR benchmark, but also showed the least degradation on general tasks. We believe this situation is related to the learning rate settings (refer to Appendix D for further discussion).
4.2 Task 2: ToolACE Function-Calling Dataset
Model Overall Acc (%, ↑) Single Turn Acc Multi Turn (%, ↑) Hallucination Measurement
Non-Live (%, ↑) Live (%, ↑) Relevance (%, ↑) Irrelevance (%, ↓)
ToolACE 58.32 87.56 76.10 7.62 83.33 88.05
LoKI($q{=}10$) 56.76 80.81 70.46 16.50 83.33 82.21
LoKI($q{=}20$) 58.16 81.02 73.26 17.75 83.33 85.73
LoKI($q{=}30$) 58.93 82.71 75.70 16.25 83.33 87.65
LoKI*($q{=}30$) 57.16 81.96 71.66 15.75 83.33 84.01
Table 3: Performance comparison on Berkeley Function Calling Leaderboard V3. ToolACE denotes the official model trained using LoRA. * indicates that we incorporated LoRA during training in this model.
Model #Params TriviaQA GSM8K Hellaswag WinoGrande HumanEval IFEval Avg(%, ↓)
Llama3.1 NaN 65.77 84.46 73.85 62.98 68.29 79.76 NaN
LoKI($q{=}10$) 188M 65.85 84.99 75.14 61.33 68.90 77.54 0.34
LoKI($q{=}20$) 376M 65.60 84.31 73.60 60.62 67.68 78.18 0.93
LoKI($q{=}30$) 563M 65.49 84.61 71.16 61.09 70.73 77.94 1.23
LoKI*($q{=}30$) 16M 64.30 83.47 70.78 62.51 70.12 78.21 1.26
ToolACE 42M 64.63 79.53 46.81 42.78 60.37 72.76 16.11
Table 4: Benchmark scores of models fine-tuned on ToolACE Function-Calling Dataset. Llama3.1 denotes Llama3.1-8B-Instruct.
Table 3 demonstrates that models trained using LoKI exhibit performance comparable to ToolACE, which was trained using LoRA on all modules. Specifically, LoKI($q{=}30$) achieves the highest overall accuracy (58.93%), surpassing the official ToolACE baseline. Notably, LoKI($q{=}20$) exhibits the strongest multi-turn reasoning capability with a 17.75% success rate, significantly outperforming ToolACE’s mere 7.62%. All LoKI variants reduce indicators of Irrelevance, which may suggest the ability of LoKI to alleviate the model’s hallucination after fine-tuning.
Table 4 further reveals LoKI’s unique capability in preserving foundational model competencies. Despite ToolACE’s 16.11% average performance drop across six benchmarks, LoKI($q{=}30$) maintains near-original capabilities with only 1.23% degradation, particularly evident in complex reasoning tasks like GSM8K (84.61% vs ToolACE’s 79.53%) and programming challenges (70.73% HumanEval accuracy vs 60.37%). This preservation consistency holds across all model scales.
Additionally, we experimented to investigate the potential of combining LoRA with LoKI (LoKI*($q{=}30$)). As anticipated, the number of trainable parameters in the model significantly decreased when integrated with LoRA, showing a reduction of 97.16% compared to LoKI($q{=}30$). Importantly, this combination did not visibly compromise LoKI’s ability to resist catastrophic forgetting. Although its performance on specialized tasks after fine-tuning is not as high as that of other models, we believe that this combination holds significant promise for future optimization.
5 Ablation Studies
In this section, we evaluate (1) the precision of KVA in quantifying knowledge vectors’ contributions to general task performance, and (2) the effectiveness of the Layer-Balanced Strategy in protecting model general task performance.
5.1 Effect of Knowledge Vector Attribution
We use Llama3.1-8B-Instruct to validate if KVA can effectively attribute the contribution of knowledge vectors to general tasks. We compare the base model with two suppression strategies (i.e., zeroing the corresponding weights):
  • S-H/L: Suppresses the top $q$knowledge vectors with the highest/lowest attribution scores (expected to significantly/minimally degrade performance if KVA correctly identifies critical vectors)
Table 5: Performance of two suppression strategies on benchmarks when $q{=}1$.
Model TriviaQA GSM8K HellaSwag WinoGrande HumanEval IFEval Avg(%, ↓)
Llama3.1 65.77 84.46 73.85 62.98 68.29 79.76 NaN
S-H($q{=}1$) 62.32 15.16 23.54 60.93 43.29 59.73 36.73
S-L($q{=}1$) 62.05 57.85 73.84 60.3 54.88 74.15 11.35
As presented in Table 5, when the top 1% of high-contribution vectors are suppressed, there is a significant performance decline, with an average degradation of 36.73%. In contrast, the performance gap between the two suppression strategies is 25.38%, which indicates significant differences and supports the accuracy of KVA.
5.2 Effect of Layer-Balanced Strategy
Next, we test the necessity of the Layer-Balanced Strategy by fine-tuning Qwen2.5-0.5B-Instruct on the LB Rerank Dataset, using two globally imbalanced schemes. We compare these results with models trained using LoKI on evaluation benchmarks.
  • G-H/L: Globally set the top $q$knowledge vectors across all layers with the highest/lowest attribution scores as trainable parameters.
As shown in Table 6, both imbalanced strategies underperform compared to LoKI: G-H shows a 39.04% average performance gap, while G-L exhibits a 30.48% average gap. The performance gaps observed in both imbalanced strategies, despite utilizing the same quota as LoKI, demonstrate the necessity of the Layer-Balanced Strategy in LoKI. We also provide the performance on the BEIR benchmark for these two models in Appendix G.
Table 6: Performance comparison of different methods.
Model TriviaQA GSM8K HellaSwag WinoGrande HumanEval IFEval Avg(%, ↓)
LoKI($q{=}10$) 20.53 37.98 21.53 47.04 24.39 33.39 8.86
G-H($q{=}10$) 11.04 29.34 10.23 17.68 25.00 26.79 39.04
G-L($q{=}10$) 9.83 30.93 8.48 43.09 24.39 27.72 30.48
6 Related Work
Catastrophic Forgetting. The challenge of catastrophic forgetting has been extensively studied in neural networks [DBLP:conf/nips/French93, kemker2018measuring], with recent works specifically analyzing its manifestation in LLMs [li_revisiting_2024, DBLP:conf/iclr/KothaSR24]. The existing methods to mitigate CF in LLMs include rehearsal-based approaches [DBLP:conf/nips/dAutumeRKY19, DBLP:conf/nips/Lopez-PazR17, DBLP:conf/iclr/ChaudhryRRE19] and regularization strategies [DBLP:journals/pami/LiH18a, DBLP:journals/corr/KirkpatrickPRVD16, deng_flattening_2021]. However, the understanding of the model’s native knowledge storage mechanism is rarely mentioned in the process of resisting CF.
Parameter-Efficient Fine-Tuning. By updating only a small subset of parameters, PEFT reduces tuning overhead: Adapters [DBLP:conf/icml/HoulsbyGJMLGAG19, DBLP:conf/iclr/HeZMBN22] insert bottleneck modules in each transformer layer; LoRA [hu2022lora] injects low-rank matrices into attention/FFN weights; prefix tuning [DBLP:conf/acl/LiL20, DBLP:journals/corr/abs-2110-07602] adds learnable "virtual tokens." Recent studies have been dedicated to improving the issue of catastrophic forgetting in PEFT methods [DBLP:conf/coling/LengX25, DBLP:journals/corr/abs-2503-02659, DBLP:conf/emnlp/WangCGXBZZGH23, DBLP:journals/corr/abs-2503-13575], but these primarily focus on the forgetting problem that arises between multiple subtasks rather than protecting the original capabilities of the model.
Knowledge Locating and Editing. In recent years, there has been a surge in techniques for modifying the knowledge of LLMs [zhang_comprehensive_2024]. For instance, ROME [meng_locating_2022] modifies the FFN weight vectors to revise one factual association, and KN [DBLP:conf/acl/DaiDHSCW22] controls the activation level of knowledge neurons to regulate the expression of relevant facts. Geva et al. [geva_transformer_2021-1, geva_dissecting_2023, geva_transformer_2022-1] conducted extensive research on how transformers intrinsically store knowledge, providing theoretical guidance for knowledge localization and editing. However, a gap remains in integrating such mechanisms into efficient fine-tuning pipelines that preserve the model’s broad pre-trained capabilities.
7 Conclusions
We present LoKI, a parameter-efficient fine-tuning framework that achieves balanced adaptation between downstream task performance and preservation of pre-trained knowledge in large language models. Our key insight stems from systematically analyzing the hierarchical knowledge storage mechanism in transformer FFN layers and developing a layer-balanced parameter selection strategy guided by integrated gradient attribution. Through experiments on retrieval and tool-use tasks, we demonstrate that LoKI achieves competitive task adaptation while significantly reducing catastrophic forgetting compared to full parameter fine-tuning and LoRA. By connecting insights from mechanistic interpretability research to practical fine-tuning goals, our study emphasizes the promising potential of this interdisciplinary field.
[…]
Appendix A Details and Discussions of KVA
A.1 Exploration of Different Settings
A.1.1 Step setting
Layer Sim. (%) Layer Sim. (%)
0 83.15 12 98.88
1 94.38 13 98.88
2 97.75 14 100.00
3 96.63 15 95.51
4 98.88 16 98.88
5 97.75 17 97.75
6 97.75 18 100.00
7 95.51 19 98.88
8 98.88 20 100.00
9 98.88 21 98.88
10 97.75 22 97.75
11 96.63 23 98.88
Total 97.42
Table 7: The similarity of each layer and the total similarity of trainable nodes obtained using two different Riemann approximation steps on Qwen2.5-0.5B-Instruction. Sim. represents similarity.
We examined the effect of the Riemann approximation step $m$ on the difference of trainable node positions determined by the Layer-Balanced Strategy ($q{=}10$). In particular, we compare the outcomes using $m{=}7$ and $m{=}20$ on Qwen2.5-0.5B-Instruct. Table 7 reports the similarity between the resulting trainable node positions under these two configurations.
To quantify the consistency of node selections between the two settings, we define similarity at both the layer level and the total network level. Let $A$ and $B$ denote the different sets of trainable node positions. Furthermore, let the network consist of $n$ layers, and $A_{i}\subset A$, $B_{i}\subset B$ denote the selected trainable node positions in the $i$-th layer under each setting.
The layer-wise similarity for the $i$-th layer is defined as:
$\text{Sim}(A_{i},B_{i})=\frac{|A_{i}\cap B_{i}|}{|A_{i}|},$ (13)
which measures the proportion of nodes in $A_{i}$ that also appear in $B_{i}$.
The overall similarity across all layers is computed as the average of layer-wise similarities:
$\text{Sim}(A,B)=\frac{1}{n}\sum_{i=1}^{n}\frac{|A_{i}\cap B_{i}|}{|A_{i}|}.$ (14)
A.1.2 Number of samples
Layer Sim. (%) Layer Sim. (%)
0 94.38 12 95.51
1 94.38 13 96.63
2 96.63 14 94.38
3 94.38 15 92.13
4 95.51 16 93.26
5 92.13 17 94.38
6 96.63 18 97.75
7 96.63 19 94.38
8 96.63 20 97.75
9 96.63 21 96.63
10 94.38 22 96.63
11 92.13 23 93.26
Total 95.13
Table 8: The similarity of each layer and the total similarity of trainable nodes obtained using two different MMLU sample selection methods on Qwen2.5-0.5B-Instruction.
We also investigated the impact of using different numbers of MMLU samples in the KVA process on the final trainable nodes. Specifically, we compared the similarity between using only the first 50 samples for each MMLU subset and using all MMLU samples. Note that we compare at $m{=}7$ and $q{=}10$. Following the definition of similarity in the previous section, we report the node similarity obtained by two methods on Qwen2.5-0.5B-Instruct and Llama3.1-8B-Instruct in Tables 8 and 9, respectively.
Layer Sim. (%) Layer Sim. (%)
0 97.07 16 97.80
1 97.07 17 98.53
2 96.82 18 97.07
3 97.31 19 96.58
4 98.29 20 97.31
5 97.07 21 98.29
6 97.07 22 97.07
7 97.56 23 97.56
8 97.07 24 98.29
9 97.80 25 98.04
10 98.04 26 97.31
11 97.56 27 98.78
12 97.80 28 97.31
13 97.56 29 98.04
14 97.80 30 98.04
15 97.56 31 96.82
Total 97.57
Table 9: The similarity of each layer and the total similarity of trainable nodes obtained using two different MMLU sample selection methods on Llama3.1-8B-Instruct.
A.2 Discussions
As mentioned in Section 3.1, the KVA process involves a one-time computational overhead, which can be considered from both the perspectives of time and computational cost. Here, we discuss potential directions for optimizing both types of overhead.
For the following two reasons, we believe that the current computational timing overhead can be significantly reduced: (a) During the KVA process of Llama3.1-8B-Instruct on a single A100 GPU, we observed that the GPU was not fully utilized, with utilization ranging from approximately 50% to 70%. We attribute this underutilization to the fixed inference batch size being set equal to step $m$. (b) The current implementation only supports single-GPU execution. We anticipate that extending the code to support multi-GPU execution will lead to a substantial reduction in computation time.
For the computational cost, as shown by our exploration of different values of m and sample sizes in this section, we believe there is still room for optimizing both the number of samples used for executing KVA on MMLU and the number of steps $m$ employed in the Riemann approximation.
In addition, we believe that exploring more suitable datasets for use in the KVA process is an appealing direction for future work.
Appendix B KVA Heatmaps
B.1 Qwen2.5-0.5B-Instruct
Figure 3: Heatmaps of top $10\%$ KVA results across all 24 FFN layers of Qwen2.5-0.5B-Instruct. The upper (red-tinted) map highlights the distribution of high-contribution node positions, while the lower (blue-tinted) map highlights the distribution of low-contribution node positions. Color intensity (log-scale) indicates the density of neurons in each category, with darker colors representing higher density.
B.2 Llama2-7B
Figure 4: Heatmaps of top $5\%$ KVA results across all 32 FFN layers of Llama2-7B, following the same visualization settings as in Fig. 3.
Appendix C Implementation of LoKI Linear
We provide the PyTorch implementation code for LoKI Linear, and readers can access the complete project code on our public GitHub repository. Based on this implementation, we performed all the LoKI-based model training in the paper. Our code has been validated for use with Llama-Factory [^2] URL omitted for double-blind review.
and supports integration with LoRA.
{minted}
[ frame=lines, linenos, breaklines, fontsize=]python import torch import torch.nn as nn
class LoKILinear(nn.Module): def __init__(self, original_linear, target_pos): super().__init__() self.out_features = original_linear.out_features self.in_features = original_linear.in_features self.active_pos = sorted(target_pos) self.frozen_pos = [ i for i in range(self.out_features) if i not in self.active_pos ]
# Parameter validation if not all(0 <= idx < self.out_features for idx in self.active_pos): raise ValueError(f"Activation indices must be within [0, self.out_features - 1]") if len(self.active_pos) != len(set(self.active_pos)): raise ValueError("Activation indices contain duplicate values") self.active = nn.Linear(self.in_features, len(self.active_pos), bias=False) self.frozen = nn.Linear(self.in_features, len(self.frozen_pos), bias=False) # Split the weight matrix W = original_linear.weight.data self.active.weight = nn.Parameter(W[self.active_pos].clone(), requires_grad=True) self.frozen.weight = nn.Parameter(W[self.frozen_pos].clone(), requires_grad=False)
# Handle bias if original_linear.bias is not None: b = original_linear.bias.data self.active_bias = nn.Parameter( b[self.active_pos].clone(), requires_grad=True ) self.frozen_bias = nn.Parameter( b[self.frozen_pos].clone(), requires_grad=False ) else: self.register_parameter("active_bias", None) self.register_parameter("frozen_bias", None)
# Pre-generate index mapping index_map = torch.empty(self.out_features, dtype=torch.long) index_map[self.active_pos] = torch.arange(len(self.active_pos)) index_map[self.frozen_pos] = torch.arange(len(self.frozen_pos)) + len( self.active_pos ) self.register_buffer("index_map", index_map)
def forward(self, x): active_out = self.active(x) # Compute active part via submodule frozen_out = self.frozen(x) # Fixed part output = torch.cat([active_out, frozen_out], dim=-1)
# Add combined bias if self.active_bias is not None: bias = torch.cat([self.active_bias, self.frozen_bias], dim=0) output += bias.unsqueeze(0).unsqueeze(0) # Broadcast bias to all batches and sequence positions
# Reorder output using pre-generated indices return output.gather( dim=-1, index=self.index_map.view(1, 1, -1).expand( output.size(0), output.size(1), -1 ), )
Appendix D Experiment Setups for the LB Reranker Dataset
Model Learning Rate Batch Size Epochs LR Scheduler WarmUp Ratio
LoKI($q{=}5$) 1.0e-5 1 1 cosine 0.01
LoKI($q{=}10$) 1.0e-5
LoKI($q{=}20$) 5.0e-6
LoKI($q{=}30$) 4.0e-6
Table 10: Training hyperparameters on LB Reranker Dataset.
On this dataset, we adopt the training parameter settings of the official full-parameter fine-tuning model, with the only modification being the learning rate. We find that when increasing the quota, appropriately lowering the learning rate helps maintain stable performance on general tasks while ensuring adaptability to downstream tasks. Table 10 shows the specific training hyperparameters for each model. Refer to the supplemental material for more detailed training hyperparameters. Due to variations in the availability of computing resources, we employed multiple hardware configurations to complete the training of these models. The configurations included the following setups: 8 RTX 4090 GPUs, 4 RTX 4090 GPUs, and 2 A100 GPUs.
We use the BEIR evaluation code [^3] URL omitted for double-blind review. provided by the dataset provider to assess our models. Specifically, we evaluate on 9 datasets from BEIR: Arguana, Dbpedia-entity, FiQA, NFCorpus, SCIDOCS, SciFact, TREC-COVID-v2, ViHealthQA, and Webis-Touche2020. For each dataset, we evaluate on a subset of the queries (the first 250).
Appendix E Experiment Setups for the ToolACE Function-Calling Dataset
Model Learning Rate Batch Size Epochs LR Scheduler WarmUp Ratio
LoKI($q{=}10$) 9.0e-06 4 3 cosine 0.1
LoKI($q{=}20$)
LoKI($q{=}30$)
Table 11: Training hyperparameters on ToolACE Function-Calling Dataset.
Model Learning Rate LoRA rank LoRA alpha
LoKI*($q{=}30$) 5.0*e-04 32 64
Table 12: Training hyperparameters on ToolACE Function-Calling Dataset when using LoRA. The parameters not mentioned in the table will continue to use the parameters in Table 11.
Apart from the learning rate and batch size, we adopt the training hyperparameters reported by the official model provider. Table 11 presents the training parameters used in our standard training approach, while Table 12 reports the hyperparameters used in the LoRA-integrated training setup. Refer to the supplemental material for complete training configuration files. All models trained on this dataset were completed using 2 A100 GPUs.
As mentioned in the main text, we use the publicly available evaluation codebase for the Berkeley Function Calling Leaderboard (BFCL) [^4] URL omitted for double-blind review..
Appendix F Benchmark Execution Details
To ensure a fair comparison of model performance, we use OpenCompass [^5] URL omitted for double-blind review. to evaluate all the models discussed in this paper, including the two base models: Llama3.1-8B-Instruct and Qwen2.5-0.5B-Instruction. The version of the benchmarks can be viewed in Table 13. Specifically, for IFEval, we calculate the average of four indicators. The scores for the remaining benchmarks are reported directly.
Dataset Version Metric Mode
winogrande 458220 accuracy gen
openai_humaneval 8e312c humaneval_pass@1 gen
gsm8k 1d7fe4 accuracy gen
triviaqa 2121ce score gen
hellaswag 6faab5 accuracy gen
IFEval 3321a3 Prompt-level-strict-accuracy gen
IFEval 3321a3 Inst-level-strict-accuracy gen
IFEval 3321a3 Prompt-level-loose-accuracy gen
IFEval 3321a3 Inst-level-loose-accuracy gen
Table 13: Specific information on all benchmarks in OpenCompass.
Appendix G Performance on the BEIR Benchmark for G-H/L
Table 14: Comparison with the full parameter fine-tuning model on BEIR.
Model MAP@1 (%) MAP@10 (%) Recall@1 (%) Recall@10 (%) NDCG @1(%) NDCG @10(%) P@1 (%) P@10 (%)
G-H($q{=}10$) -1.0 -2.8 -1.0 -1.6 0.4 -1.2 0.0 -1.0
G-L($q{=}10$) -4.5 -4.8 -4.5 -2.8 -2.6 -3.0 -2.6 -3.0
Table shows the percentage difference in performance compared to the full parameter fine-tuning model across standard retrieval metrics.
Appendix H Limitations and Future Works
[…]
Appendix I Broader Impacts
Positive Societal Impacts. LoKI enables LLMs to be adapted to new tasks with minimal loss of pre-trained general knowledge. By preserving general capabilities, LoKI promotes the development of multi-domain models that retain versatility across a wide range of tasks, potentially reducing the need to maintain and manage a large number of specialized models. This can foster more unified and maintainable AI systems with consistent behavior across applications.
Negative Societal Impacts and Risks. As with any technique that enhances the efficiency and accessibility of LLM customization, LoKI may inadvertently lower the barrier for misuse. Specifically, malicious actors could leverage LoKI to implant biased, harmful, or deceptive capabilities into powerful language models while preserving general fluency, potentially making such manipulations harder to detect. Additionally, while LoKI protects general knowledge, the process of knowledge implantation could still result in subtle shifts in behavior that are difficult to audit.
Moreover, the reliance on pretrained LLMs—often trained on opaque or uncurated datasets—means that any biases present in the original model are likely to persist or be reinforced during adaptation. LoKI does not directly address these foundational concerns, and care must be taken to ensure its application aligns with ethical and fairness principles.
Mitigation Strategies. To mitigate risks, we encourage the use of LoKI in conjunction with auditing tools, attribution techniques, and interpretability methods to ensure transparency in model modifications. Furthermore, deployment should be accompanied by rigorous evaluation on fairness, robustness, and safety metrics, especially when the adapted models are intended for high-stakes applications.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: In the abstract and introduction we introduce the current situation of PEFT and CF. We claim that our work provides a PEFT approach to alleviate CF while maintaining downstream task performance.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations of our work in Appendix H.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We include theoretical justifications for our method in Section 3.1.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We explain our experiments in Section 4 and provide the complete set of hyper-parameters in Appendix D and Appendix E.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide a zip file containing the code for our experiments, and the GitHub repository will be made public.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so "No" is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide information about the training details in Appendix D and Appendix E.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We do not provide error bars, but we explained the hyperparameter selection criteria for all experiments in Appendix D and Appendix E. We only adjusted the learning rate to suit various quotas and documented the best results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: All these information can be found in Appendix D and Appendix E.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in this paper conforms, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss broader impacts in Appendix J.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We do not release any models.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We work with open-source models and datasets that are publicly available, and we cited them properly.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide code and instructions on how to run the code to reproduce our results.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: The method proposed in this paper mainly focuses on LLMs as the execution object.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.